forty ounces of a 30% gold alloy are mixed with 54 ounces of a 20% alooy. Find the concentration of teh resulting gold alloy.
Answer by Alan3354(30993)

forty ounces of a 30% gold alloy are mixed with 54 ounces of a 20% alooy. Find the concentration of teh resulting gold alloy.
--------------------------
40 oz x 0.3 = 12 oz of gold
54 oz x 0.2 = 10.8 oz of gold
12 + 10.8 = 22.8 oz of gold
40 + 54 = 94 oz total
22.8/94 = 0.24255...
= 24.26%
-----------------------------
The answer seems reasonable, as it's between the 2 percentages, and closer to the one that most of was used.
Always check your answers for reasonableness. If a mistake had been made, and the answer came out as 48%, or 12%, you should recognize that something is wrong.